OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (lgbasallote):

use (4.6) and (-1,-4) as slope...then use (-1,1) for x1 and y1 in \[\huge y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] does that help?

OpenStudy (anonymous):

no really what would the equation be if you substitute the number in?

OpenStudy (lgbasallote):

kinda like \[\huge y - 1 = \frac{-4 - 6}{-1 - 4} [x - (-1)]\]

OpenStudy (anonymous):

so y-1=-2[x+1]???

OpenStudy (lgbasallote):

close...but not -2

OpenStudy (anonymous):

oh positive 2?

OpenStudy (lgbasallote):

yup..now simplify the equation

OpenStudy (anonymous):

thats what i need your help.. how do i do that from here?

OpenStudy (lgbasallote):

wait...i just realized it was just asking for slope o.O well the slope is just 2...

OpenStudy (anonymous):

oh i thought so! haha thank you very much

OpenStudy (lgbasallote):

welcome ^_^